OpenStudy (anonymous):

A ray of light goes from air into the substances shown in the table at an angle of 30 degrees. Which substance will bend the light the most? water , flint glass, diamond, quartz ?

OpenStudy (anonymous):



rvc (rvc):

denser to rarer medium the ray with be away from the normal

OpenStudy (michele_laino):

here we have to apply the law of Snell: \[\large {n_1}\sin {\theta _1} = {n_2}\sin {\theta _2}\]

OpenStudy (anonymous):

ok!

rvc (rvc):

yep

OpenStudy (michele_laino):